# A Reference Model and Patterns for Production Event Data Enrichment

### v1

@@ 1 Introduction
We extract the patterns from observations on data sets from four different industrial production environments. To describe and formalise the patterns, we defined a reference model. The reference model combines the well-known standard ISA-95 [31] and the Event Knowledge Graph formalism [14, 15]. The benefit of using this reference model is that it provides both a uniform terminology to describe the patterns and a way to automatically apply the patterns to the data stored according to the reference model. Subsequently, we define each of the patterns in terms of the reference model and motivate them by providing use cases. We also show in which of the four different industrial production environments studied the patterns are observed.
@@ 3 Methodology
We identified the patterns based on observations about existing data sets and subsequently formalised them using a reference model that was constructed based on existing standards. This section describes the way in which the patterns are derived from the data sets and the way in which the reference model was constructed from the existing standards.
@@ 4.1 Reference Model
In this section, we describe the main concepts of the reference model. An overview of our model can be found in Figure 2.
@@ 5 Implementation and evaluation
The patterns are implemented as SPARQL [38] query templates [^5] https://github.com/gitmpje/production-trace-patterns/tree/main/sparql. Those queries can be used to extract information from the event data and enrich it. The repository also contains a dummy data set and a script to apply the patterns for a selection of the use cases.
As can be seen in Table 3, most of the use cases are observed in multiple data sets. If we aggregate the results over the patterns, we can see that most of the patterns apply to at least three data sets, except for Pattern 4, 6, and 9, which appear to be more specific for a certain data set.
@@ 6 Conclusion and Discussion
The digital transformation has led to an increase in data collection across various systems. When data from these systems are combined, they yield new and valuable insights. Knowledge about the context and data of the process is required to extract this hidden information. Therefore, this information is often extracted ad hoc during data pre-processing. In examining event data sets from four production environments, we identified recurring themes in the information of interest. This paper introduces production trace patterns, generic patterns to uncover hidden information. These patterns are defined using a reference model and are illustrated through industry use cases. To our knowledge, this is the first work in which patterns are described for deriving hidden information from event data sets from manufacturing processes.
We recognise that restricting our study to data sets and use cases in the (discrete) manufacturing field is a significant limitation of this work. This is because EKGs, which underpin our reference model, are primarily used for multi-dimensional event data from manufacturing processes. Nevertheless, this work serves as an initial step, and we anticipate the applicability of the abstract patterns in other fields. For example, Pattern 1 could be used in a data centre, where a computing Resource runs a programme to handle a job (ProductionEntity), with alarm Events triggered by exceptions.
We do not assert that the patterns gathered here are complete. Each manufacturing setting is unique, necessitating distinct patterns to uncover hidden information. Thus, there is room for adapting and expanding patterns across diverse manufacturing environments, not limited to discrete manufacturing. Collecting patterns is a step forward towards systematically enriching event data. We hope the community will embrace and apply these patterns, growing the collection over time. The broader the pattern set, the more engineers and practitioners can benefit from new insights in their event data.

### v2

@@ 1 Introduction
We derive the patterns from observations about data sets from four different industrial production environments. To describe and formalise the patterns, we define a reference model. The reference model combines the well-known standard ISA-95 [31] and the Event Knowledge Graph formalism [14, 15]. The benefit of using this reference model is that it provides both a uniform terminology to describe the patterns and serves as a basis for extracting information from databases that are structured according to those standards. Subsequently, we define each of the patterns in terms of the reference model and motivate them by providing use cases. We also show in which of the four different industrial production environments that we studied the patterns are observed.
@@ 3 Methodology
We identified the patterns based on observations about existing data sets and subsequently formalised them using a reference model that was constructed based on existing standards. This section first describes the way in which the reference model was constructed from existing standards and then the way in which the patterns are derived from the data sets.
@@ 4.1 Reference Model
In this section, we describe the main concepts of the reference model. The concepts are based on existing definitions and standards. Therefore, the model also serves as a basis for extracting information based on the patterns when databases are structured according to those concepts. An overview of our model can be found in Figure 2.
@@ 5 Implementation and evaluation
The patterns are implemented as SPARQL [38] query templates [^5] https://github.com/gitmpje/production-trace-patterns/tree/main/sparql. Those queries can be used directly in systems that are built on the standards on which the reference model is based. This makes the patterns practically applicable for extracting information from event data and enriching it. The repository also contains a dummy data set and a script to apply the patterns for a selection of the use cases, showing that the patterns correctly extract information.
In addition, we evaluated the applicability of the patterns by recording their appearance in the data sets. As can be seen in Table 3, most of the use cases are observed in multiple data sets. If we aggregate the results over the patterns, we can see that most of the patterns apply to at least three data sets, except for Pattern 4, 6, and 9.
A use case for Pattern 4 is the downtime of production resources, which is also relevant in other settings, but captured by different types of events. In these settings, Pattern 3 can be used to derive the information. Pattern 6 and 9 are more specific for production processes in which large volumes of products are processed in batches, which we only observed in one of the data sets.
In general, the patterns are broadly applicable to a range of manufacturing settings. This also implies that the concepts from the reference model are well represented in the data sets.
@@ 6 Conclusion and Discussion
The digital transformation has led to an increase in data collection across various systems. When data from these systems are combined, they yield new and valuable insights. Knowledge about the context and data of the process is required to extract this hidden information. Therefore, this information is often extracted ad hoc during data pre-processing. In examining event data sets from four production environments, we identified recurring themes in the information of interest. This paper introduces generic production trace patterns to uncover hidden information. These patterns are defined using a reference model. The reference model is based on existing definitions and standards and, therefore, also serves as a basis for extracting information in settings where data are structured according to those standards. We evaluated the patterns by showing their appearance data sets from practice and demonstrate that they correctly extract information by implementing them in SPARQL. To our knowledge, this is the first work in which patterns are described for deriving hidden information from event data sets from manufacturing processes.
We recognise that restricting our study to data sets and use cases in the (discrete) manufacturing field is a significant limitation of this work. This is because EKGs, which underpin our reference model, are primarily used for multi-dimensional event data from manufacturing processes. Nevertheless, this work serves as an initial effort and we anticipate the applicability of the abstract patterns in other fields. For example, Pattern 1 could be used in a data centre, where a computing Resource runs a programme to handle a job (ProductionEntity), with alarm Events triggered by exceptions.
We do not claim that the patterns gathered here are complete. Each manufacturing setting is unique, necessitating distinct patterns to uncover hidden information. Thus, there is room for adapting and expanding patterns in diverse manufacturing environments, not limited to discrete manufacturing. Nevertheless, the collection of patterns and their implementation in SPARQL using a reference model based on standards makes them practically applicable and is a step towards systematically enriching event data. We hope that the community will embrace and apply these patterns, growing the collection over time. The broader the pattern set, the more practitioners can benefit from new insights in their event data.